Ok My issues are 1/2 solved.
I fixed the Top Popper Lock. I traced the switch wires with the matirx chart and found that the GRN-GRY wire on the Left Outlane (60) was touching the base of one of the new LEDs I put in. I moved that away and now the Top Popper Lock is working again 100%. Woot!
Now just got to work on why my Right Troll Guards and not working correctly. No response to Edge test on any of the 3. Does not set off Troll when they are dropped. Once ball drops to bottom, out of play they do reset back up physically. As mentioned above just finished cleaning the sensors, no go. Possible I need a new board for them?
Left Troll guards and Troll works 100% and since they are on the same row switch wise doubt it's a short like my last problem.
*So might be possible this issue also shorted out the right drop targets board as the Right Drop Target Top switch is on the same wire as the Left Outlane that had the short.
Right Troll Guard issues were a wire that had come off the bottom of the molex on the Left guards. It jumps over the the right side drop targets so it was not getting power. Fixed that wire in the Molex and working 100%.
Thanks for all the troubleshooting help everyone. I learned a bunch from this so I'm stoked about that.